OpenStudy (anonymous):

Factor the difference of two cubes. x^3n-729

OpenStudy (anonymous):

Two cubes can be factored as such: \[ a^3+b^3=(a+b)\left(a^2-ab+b^2\right) \]Plug in \(-b\) for a negative cube.

OpenStudy (anonymous):

(x^n)^3-(9)^3

OpenStudy (anonymous):

DOES IT HELP?

OpenStudy (anonymous):

yes so what is my next step?

OpenStudy (anonymous):

or is that the answer

OpenStudy (anonymous):

HInt: a^3-b^3=(a-b)(a^2+2ab+b^2)

OpenStudy (anonymous):

so my answer is going to be : (a^2+2(3)(3)+3^2)?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

(x^n)^3-(9)^3 (x^n -9) (x^2n + 18x^n + 81) (x^n -9) {(x^n)^2+2 * x^n * 9 + 9^2} (x^n -9) (x^n+9) (x^n +9)

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

so all 3 of them on the bottom are my answer?

OpenStudy (anonymous):

YEP

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@sauravshakya, you have your base equation wrong, it is *not* \(a^2+2ab+b^2\), it is \(a^2+ab+b^2\)!

OpenStudy (anonymous):

Oh ya........ thanx @LolWolf

OpenStudy (anonymous):

@KE8717504 u there